OpenStudy (anonymous):

What operation is necessary to solve the equation that represents the statement: two point four is 8 multiplied by what number? Explain your answer.

OpenStudy (anonymous):

By two point four you mean 2.4?

OpenStudy (anonymous):

@icat21 Yes :)

sammixboo (sammixboo):

2.4 = 8*n ^That is what you are basically what you are trying to solve... You have to first divide 8 on both sides.

sammixboo (sammixboo):

I am not sure how to explain it but all you have to do is divide 8 on both sides of the equation

OpenStudy (anonymous):

Can you give me an example?

OpenStudy (anonymous):

2.4=8*n where n is a number. What you need to do is divide from both sides. 2.4/8=n n=0.3

sammixboo (sammixboo):

Here - 2.4 = 8*n 2.4/8 = 8*n/8 (in this step the 8's get cancelled out) 0.3 = n

OpenStudy (anonymous):

Thank you :)